# Characterizing randomness in parameterized quantum circuits through expressibility and average entanglement

### v1

@@ I Introduction
In this work, we discuss the expressibility and average multipartite entanglement of parameterized quantum circuits with restrictions on the qubits connectivities. Applying the expressibility [26] and the average entanglement for the Scott quantifier of order $1$ and $2$ [44], we discuss the consequences of the restrictions on the generation of quantum states. We also show that, in fact, these quantities are connected and the topology of the quantum architecture has consequences for the observed features. PQCs generating random states with average and standard deviation of entanglement closer to the observed for the ensemble of uniformly distributed random unitaries have a steeper increase in expressibility. The paper is structured as follows. In Section II we introduce the PQCs investigated in this work and the motivation for the choices. In Section III we define the quantifiers applied for the investigation, i.e., the expressibility and the average Scott quantifier. In Section IV, we discuss the results for the proposed PQCs. Finally, the conclusions are presented in Sec. V.
@@ II.1 Circuit Architectures and Ansätze
We chose local operations RX and RY due to the fact that they are going to generate random $1$ qubit states over the complete surface of the Bloch sphere sampling the circuit parameters uniformly, when starting with a |$\ket{0}$ state. One of the drawbacks of this choice is that the random $1$ qubit states will present a bias towards two of the axis of the sphere, however the states generated effectively cover the space [26]. Table 1 presents the number of CNOT gates and total number of gates in both circuits Ansatz $1$ and $2$. Comparing the different topologies, as a function of the number of qubits, $n$, and number of layers, $l$. The total number of parameters appears isolated inside brackets in the equations of the total number of gates, so this number can be easily known looking at the second term inside brackets. For example, for the Linear topology and Ansatz $2$, there are $\left[(n-1)+4n\right]l$ gates, being $(n-1)l$ CNOTs and $(4n)l$ parameterized gates.
@@ III.1 Expressibility
This measure over the unitary group is called Haar measure and present interesting properties as invariance considering elements of the group [59]. However, build a faithful replication of the Haar measure can be hard and costly, requiring circuits with gate complexity that scales exponentially with the dimension of the system [30]. Also, only part of the properties of the Haar measure are needed to determine integrals of polynomials of restricted order or mean values of operators, motivating the definition of $t-$designs [60, 61, 62]. The $t-$designs are ensembles of unitaries that can replicate parts of the characteristics of the Haar measure over the unitary group, with different plausible equivalent definitions [59]. One possible definition is in terms of the polynomials one might want to calculate integrals over the group manifold.
@@ III.2 Scott entanglement measure and average entanglement
The different values of $m$ in the Scott entanglement measure will explore different characteristics of multipartite entanglement and will define different entanglement measures. For example, there are states for which the MW, $m=1$, will present the same entanglement values, even though they have very different characteristics. We can cite the states of $6$ qubits, |$\ket{GHZ_{6}}=(\ket{0}^{\otimes 6}+\ket{1}^{\otimes 6})/2$ and |$\ket{EPR_{6}}=\ket{\Phi^{+}}^{\otimes 3}$, being |$\ket{\Phi^{+}}=(\ket{00}+\ket{11})/\sqrt{2}$, that will have $Q_{1}(\ket{GHZ_{6}})=Q_{1}(\ket{EPR_{6}})=1$. These states have very different entanglement characteristics, being the |$\ket{GHZ_{6}}$ sometimes considered a state with maximum multipartite entanglement, while the |$\ket{EPR_{6}}$ is a type of biseparable state [68, 69]. Choosing $m=2$, this degeneracy is broken and we have $Q_{2}(\ket{GHZ_{6}})=2/3$ and $Q_{2}(\ket{EPR_{6}})=4/5$. Therefore, in this work we compute the average entanglement considering the measures $S_{1}$ and $S_{2}$, so we can understand if the results can be replicated with both entanglement measures.
@@ Acknowledgements.
This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brazil (CAPES) – Finance Code 001. I.M. acknowledges financial support from São Paulo Research Foundation - FAPESP (Grant No. 2022/08786-2 and No. 2023/14488-7). P.C.A. acknowledges financial support from Conselho Nacional de Desenvolvimento Científico e Tecnológico - Brazil (CNPq - Grant No. 160851/2021-1). D.O.S.P acknowledges the support by the Brazilian funding agencies CNPq (Grant No. 304891/2022-3), FAPESP (Grant No. 2017/03727-0) and the Brazilian National Institute of Science and Technology of Quantum Information (INCT/IQ).

### v2

@@ I Introduction
In this work, we discuss the expressibility and average multipartite entanglement of parameterized quantum circuits with restrictions on the qubits connectivities. Applying the expressibility [25] and the average entanglement for the Scott quantifier of order $1$ and $2$ [43], we discuss the consequences of the restrictions on the generation of quantum states. We also show that, in fact, these quantities are connected and the topology of the quantum architecture has consequences for the observed features. PQCs generating random states with average and standard deviation of entanglement closer to the observed for the ensemble of uniformly distributed random unitaries have a steeper increase in expressibility. The paper is structured as follows. In Section II we introduce the PQCs investigated in this work and the motivation for the choices. In Section III we define the quantifiers applied for the investigation, i.e., the expressibility and the average Scott quantifier. In Section IV, we discuss the results for the proposed PQCs. Finally, the conclusions are presented in Sec. V. Table LABEL:tab:summary summarizes the concepts and results considered in this work.
@@ II.1 Circuit Architectures and Ansätze
We chose local operations RX and RY due to the fact that they are going to generate random $1$ qubit states over the complete surface of the Bloch sphere sampling the circuit parameters uniformly, when starting with a |$\ket{0}$ state. One of the drawbacks of this choice is that the random $1$ qubit states will present a bias towards two of the axis of the sphere, however the states generated effectively cover the space [25]. There are other possible approaches to explore the $1-$qubit Hilbert space uniformly, however they would require different probability distributions or gate configurations. Here, we chose to follow the choice provided by the original paper where the expressibility and entangling capability were first introduced, considering the uniform distribution in a fixed interval to not infer any bias over the parameters, and RX and RY gates to be faithful to our goal of applying parameterized rotations. Table 1 presents the number of CNOT gates and total number of gates in both circuits Ansatz $1$ and $2$. Comparing the different topologies, as a function of the number of qubits, $n$, and number of layers, $l$. The total number of parameters appears isolated inside brackets in the equations of the total number of gates, so this number can be easily known looking at the second term inside brackets. For example, for the Linear topology and Ansatz $2$, there are $\left[(n-1)+4n\right]l$ gates, being $(n-1)l$ CNOTs and $(4n)l$ parameterized gates.
@@ III.1 Expressibility
This measure over the unitary group is called Haar measure and presents interesting properties as invariance considering elements of the group [58]. However, building a faithful replication of the Haar measure can be hard and costly, requiring circuits with gate complexity that scales exponentially with the dimension of the system [29]. Also, only part of the properties of the Haar measure are needed to determine integrals of polynomials of restricted order or mean values of operators, motivating the definition of $t-$designs [59, 60, 61]. The $t-$designs are ensembles of unitaries that can replicate parts of the characteristics of the Haar measure over the unitary group, with different plausible equivalent definitions [58]. One possible definition is in terms of the polynomials one might want to calculate integrals over the group manifold.
@@ III.2 Scott entanglement measure and average entanglement
The different values of $m$ in the Scott entanglement measure will explore different characteristics of multipartite entanglement and will define different entanglement measures. For example, there are states for which the MW, $m=1$, will present the same entanglement values, even though they have very different characteristics. We can cite the states of $6$ qubits, |$\ket{GHZ_{6}}=(\ket{0}^{\otimes 6}+\ket{1}^{\otimes 6})/\sqrt{2}$ and |$\ket{EPR_{6}}=\ket{\Phi^{+}}^{\otimes 3}$, being |$\ket{\Phi^{+}}=(\ket{00}+\ket{11})/\sqrt{2}$, that will have $Q_{1}(\ket{GHZ_{6}})=Q_{1}(\ket{EPR_{6}})=1$. These states have very different entanglement characteristics, being the |$\ket{GHZ_{6}}$ sometimes considered a state with maximum multipartite entanglement, while the |$\ket{EPR_{6}}$ is a type of biseparable state [67, 68]. Choosing $m=2$, this degeneracy is broken and we have $Q_{2}(\ket{GHZ_{6}})=2/3$ and $Q_{2}(\ket{EPR_{6}})=4/5$. Therefore, in this work we compute the average entanglement considering the measures $S_{1}$ and $S_{2}$, so we can understand if the results can be replicated with both entanglement measures.
@@ IV.2 Ansatz 2
IV.3 Potential limitations of the analysis
In our work we decided to follow a particular reasoning to choose the particular gates performed and connectivities between qubits, however there are biases and limitations of these choices.
We chose circuits only with local parameterized gates and CNOTs for connections between qubits. This allowed us to isolate the influences of either connectivity or parametrization, which was important to understand how ansätze $1$ and $2$ provide different answers to the generation of entanglement and expressibility. On the contrary, parameterized two qubits gates would be an interesting choice if the objective is to maximize the expressibility with circuits of low depth, as exemplified by Ref. [25], whose circuits only with CNOTs as two qubits gates present a slower decay of the $D_{KL}$. On top of that, CNOT gates are unidirectional asymmetric gates, which only allow to think about topology in a non-directional graph sense. Other promising analyses, more representative and more general, would be to consider CNOT gates in both the directions of each edge of the graph or the application of symmetric Controlled-Z gates, as done in Refs. [26, 27].
Inspired by Ref. [25], we considered the uniform distribution for each parameter to understand an unbiased behavior of the circuits with respect to the parameter choices: each parameter is as probable as any other within the interval. This is a reasonable choice when the lack of a cost function and a specific problem is considered, however it does not illustrate the general behavior of the circuit. Changing the distribution would affect the behavior observed for the expressibility and average entanglement, a priori. This is well pictured by the example given in the discussions, where the uniform distribution of one qubit states can be replicated considering RX and RY rotations if one considers the $2-$sphere inspired probability distribution. The same is not possible only with the two rotations and the uniform distribution. Furthermore, different types of chaotic dynamics emerge from different gates choices, as discussed in Refs. [26, 27], where either the eigenvalues distributions of random matrices following the Gaussian Orthogonal Ensemble or the Gaussian Unitary Ensemble are generated by the related circuit structures.
When it comes to the real hardware of quantum computers usually not only one topology is available, being the circuits built with compositions of the topologies presented in this work [52, 23]. For example, in the IBM_Ithaca quantum computer [52], it is possible to build circuits with $12$ qubits in a Ring topology, $4$ qubits in a Star topology, many configurations of qubits in a Linear topology, and, of course, any possibility of No Connections circuits. This way, we analysed only subgraphs of the more general graph that is going to be the circuit. One work that considers this kind of possibility and discusses the consequences of a combination of the All-to-all and Ring cases is Ref. [77]. From another viewpoint, the topologies will be implementable in different settings depending on the quantum computer available. For instance, ion trap quantum computers have a large freedom on the implementation of connections [54], being able to perform even the most complex All-to-all topology. Nevertheless, not every hardware of this kind will be able to implement the gates we chose in this work natively. For instance, the ion trap discussed in Ref. [78], which implements two qubit operations between every possible pair of qubits, can only implement RXX, RYY, and RZZ  [^3] These gates are defined as RII$(\theta):=\exp\left(-i\frac{\theta}{2}\sigma_{I}\otimes\sigma_{I}\right)$, being $\sigma_{I}$ the Pauli matrices with $I=X,Y,Z$. and one qubit rotations natively. A composition of these gates would be necessary to implement the CNOT, increasing the circuit size.
@@ V Conclusions
Since the introduction of the expressibility and average entanglement quantifiers as defined in Ref. [25], no other work investigated the connections between the entanglement standard deviation of the circuits at only $1$ layer and the expressibility increase as a function of circuit depth. The average entanglement growth and standard deviation decay as a function of the number of layers together with the randomness increase is a known characteristic of random circuits, and has been appreciated in many important works [42, 43, 24, 73, 28, 76, 41]. Still, the efforts to understand randomness and entanglement generation in parameterized quantum circuits are even more recent [80, 81, 28]. Leading results have investigated how circuits with this kind of structure will be able to generate $t-$designs [80, 76], will have an increase of entanglement that presents a universal increase independent on the number of qubits [28], a result also observed here, and finally how this randomness generation is influenced by the introduction of noise [81]. Recently, how the randomness in PQCs can be compared to other kinds of random circuits [82] has been explored, including analysis applying the majorization criterion for complexity quantification. Parameterized quantum circuits are one of the essential ingredients of Variational Quantum Algorithms, and the efforts to extend and understand their possibilities are vast, including results for quantum batteries and thermodynamics [83, 84], reproduction of search algorithms in the NISQ era [85], and Quantum Machine Learning [13, 86]. Ref. [9] provides an extensive review of the many possible applications of PQCs in VQAs.
@@
The results here place a connection between entanglement generation and complexity of circuits as quantified by the expressibility. A natural extension of this discussion is how these can be connected to notions of information scrambling, where local information in the input system cannot be accessed locally in the output [87]. Whether or not the same behavior is observed, i.e., circuits that are more scrambling after one layer present a steeper increase of expressibility, the relations between entanglement and information scrambling in this setting would benefit from the insights provided. From another viewpoint, the findings could be analysed in comparison with the performance of specific cases of VQA applications, for instance the Variational Quantum Eigensolver search for ground states, to understand which are the consequences to trainability. A result already observed is that circuits whose topology is completely different from the connections observed in the Hamiltonian will have a worse performance when compared to circuits with topology following the Hamiltonian structure [20]. However, how this connects to the average entanglement and expressibility is still not clear. Examples of this approach are presented in Refs. [26, 27]. As discussed in Sec. IV.3, the circuits here do not include all potential circuits appearing in quantum computers. In this sense, another promising extension of this work would be to understand how the use of different topologies in the same quantum circuit will affect the behavior of the expressibility and entanglement generation. Finally, bounds on the entanglement of quantum states were already applied to determine the possible classical simulability of quantum dynamics [88], and the complexity of quantum circuits was shown to be bounded below by their entangling power [89]. In this sense, the observed connection between expressibility and average entanglement at only one layer can be explored to understand circuit complexity growth in a more fundamental sense, and is an interesting research project for future works.
@@ Acknowledgements.
This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brazil (CAPES) – Finance Code 001. G.I.C. acknowledges the financial support of the Research Council of Finland through the Finnish Quantum Flagship project (358878, UH). I.M. acknowledges financial support from São Paulo Research Foundation - FAPESP (Grant No. 2022/08786-2 and No. 2023/14488-7). P.C.A. acknowledges financial support from Conselho Nacional de Desenvolvimento Científico e Tecnológico - Brazil (CNPq - Grant No. 160851/2021-1). D.O.S.P acknowledges the support by the Brazilian funding agencies CNPq (Grants No. 304891/2022-3 and No. 402074/2023-8), FAPESP (Grant No. 2017/03727-0) and the Brazilian National Institute of Science and Technology of Quantum Information (INCT/IQ).